# Reducing the Cost of Energy Differences in Variational Monte Carlo with Spotlight Sampling

### v1

@@ II.6 Initialization
We begin by placing electrons according to a Lewis structure so that the molecule is roughly spatially and spin neutral. The position of each electron (except for those in the carbon and oxygen cores) is then randomized within a 1 Bohr box centered on its Lewis placement. After separating the nuclei into fragments — in our tests, each carbon and any hydrogens or O-H groups bonded to it form a fragment, as do individual H2 molecules — we then group the initial electron positions by fragment so that each fragment gets the $n_{e}$ nearest electrons, where $n_{e}$ is the fragment’s nuclear charge. We burn in the electron positions using a “round-robin” approach in which we loop over each fragment and, for each one, unfreeze and sample its electrons and those of its two nearest neighbors in each direction. At the beginning of each fragment’s burn in, and once after they have all been burned in, we re-invert $M$, for a total of $n_{\mathrm{fragments}}+1$ large matrix inversions. In the systems we test, these inversions together cost much less than the initial Hartree-Fock calculation. After the round-robin burn in, we reassign electrons to fragments, again taking the closest so many so that the fragment is neutral. At this point, we have our initial positions from which we separately launch the Markov chains that evaluate each fragment’s multipole and then continue on to evaluate each fragment’s energy using the approximate Hamiltonian in Eq. (8).
III Results
III.1 Computational Details
@@
We performed all quantum chemistry calculations for this project using the PYSCF package.Sun (2015); Sun et al. (2018, 2020) We optimized all geometries at the MP2/6-31G level of theory unless otherwise specified. For our Hartree-Fock references, we used restricted Slater determinants with localized Foster-Boys orbitalsFoster and Boys (1960) using the 6-31G(d) basis set. To ensure electron-nuclear cuspsKato (1957) were satisfied for the VMC sampling, we employed a recently developed cusping algorithmQuady, Bumann, and Neuscamman (2025) in which all AOs are explicitly cusped at each nucleus. We used a simple two-body Jastrow factorFoulkes et al. (2001); Ceperley (1978); Ceperley and Alder (1980) to satisfy the electron-electron cusps (A parameter = 0.03). Regular and spotlight sampling VMC calculations were performed using our own pilot code.
@@
III.2 Impact of the Buffer Region
@@
Figure 4: Each fragment’s average uncertainty contribution to the O-H bond stretching energy in 1-octonol. Each fragment contains one C atom and any H or O-H bonded to it.
III.3 Cost Scaling
In investigating whether observed computational cost scaling matches the expectations from our analysis above, let us begin with a simple test in which we take a fixed number of samples for each fragment. As our implementation does not exploit orbital locality, we expect each sample to carry an $O(N)$ cost, which, when combined with the fact that there are $O(N)$ fragments, leads us to expect $O(N^{2})$ cost growth overall. We test this theory in a chain of hydrogen dimers (bond length = 1.393 Bohr) arranged in a line, in which each H2 molecule’s center is placed 4.03 Bohr away from the previous center, and all bond axes are arranged at an angle of 7.12 degrees relative to the line passing through the centers. Using each H2 as its own fragment and taking 110 samples per fragment, we see in Figure 3 that the spotlight sampling cost does indeed grow as $O(N^{2})$. While our current code cannot test the reduction to $O(N)$ that would result from local orbitals, that reduction is well established, Bienvenu et al. (2022); Feldt and Assaraf (2021) and so we instead turn to the question of how effectively costs can be reduced by taking fewer samples in fragments farther away from the local chemical change in question.
[…]
To address this question, we first perform tests of how much uncertainty each fragment contributes to a bond stretching energy when we take a fixed number of samples per fragment. We perform our first test in 1-octonol, where we perform 150 runs on each fragment, in which each run uses 720,000 samples. For each run in each fragment, we use a blocking analysis to evaluate the uncertainty of that fragment’s contribution to the bond stretching energy in that run. The averages of these uncertainties over all 150 runs are plotted in Figure 4, where we see that the uncertainties decrease as one moves away from the CH2OH fragment (fragment 1) that contains the bond stretch. In fact, these uncertainties decay fast enough that they are bounded by an $a/l^{2}$ function, with $a=0.01$ Hartree and $l$ the fragment number. As discussed in Section II.1, this is a highly desirable bound, as it suggests that decreasing the sample size for farther fragments should be highly effective.
@@ III.3 Cost Scaling
[…]
@@
In Figure 6, we see that this test of spotlight sampling achieves near-linear scaling with system size, with spotlight’s cost lower than standard correlated sampling once there are about 100 electrons. The linear scaling is expected, as the vast majority of the samples are in the first three fragments, making the total number of samples nearly independent of system size. Indeed, the linear cost growth comes instead from our code’s $O(N)$ cost per sample. The small quadratic term in the cost growth that begins to be visible in the spotlight timing data in the largest system is due to the fact that we used fixed-size 4,000 sample blocks even for the farthest fragments, and could be eliminated by tapering this sample size further. It is important to re-emphasize that this crossover with standard sampling has been achieved without exploiting local orbitals. Were one to do so, the crossover would occur even earlier, and the cost growth of spotlight sampling could drop to sub-linear, assuming local Jastrows and fast multipole methods were also employed. Although each spotlight prediction agreed with standard sampling within error bars, this system is somewhat artificial, and so we now turn to a series of alcohols as a more realistic test of the method’s accuracy.
III.4 Accuracy
@@
To test whether our approximations to the Hamiltonian and probability distribution produce meaningful errors in predicted energy differences, we apply the spotlight sampling approach to predict the energetic cost of a 0.2 Bohr O-H bond stretch in a series of alcohols: 1-butanol through 1-octanol. As seen in Figure 7, spotlight sampling reproduces the predictions of standard correlated sampling in all cases. We note that this accuracy comes despite the fact that, in octonol, the Markov chain that estimates the contribution from the fragment containing the stretched bond froze more than half of the electrons.
@@
Figure 7: Predictions for the energy needed to stretch an O-H bond in the alcohols 1-butanol through 1-octonol.
IV Conclusion
@@
This study has explored the efficacy of a spotlight sampling approach to predicting the energy differences associated with local chemical changes in VMC. In this approach, standard correlated sampling is approximated by using a series of separate Markov chains, one for each fragment in the molecule, in which only the electrons near that fragment move. Our theoretical analysis suggests that the cost of the approach can be made linear with system size when employing nonlocal orbitals, and that this cost should become sub-linear when local orbitals, local Jastrows, and fast multipole methods are used. Timing calculations show that the overall cost grows quadratically when we take a fixed number of samples per spotlighted fragment, and that this cost growth becomes essentially linear, as expected, once we exploit the fact that per-fragment uncertainty decays rapidly for fragments farther from the local change. To maintain accuracy despite the biases inherent to freezing most electrons’ positions, we have developed an approximate fragment Hamiltonian that employs multipoles for long range interactions. Tests on a series of alcohols show that this approach can reproduce the predictions of standard sampling within statistical uncertainty.
V Appendix
V.1 Correlated Sampling with the Space-Warp Coordinate Transformation
@@
We employ the correlated sampling Lowther and Coldwell (1980) technique known as the space-warp coordinate transformation Umrigar (1989) to reduce the statistical uncertainty of simple bond stretches (e.g., O-H stretch in 1-n-ol, n = butan-, pentan-, hexan-,…). The relative energy difference for such a stretch, $E_{s}-E$, is rewritten to be sampled from the unstretched $\Psi$ reference:
@@
$J(R_{i})$ is the Jacobian of the space-warp coordinate transformation, Umrigar (1989); Filippi and Umrigar (2000) and
@@
V.2 Kinetic Energy Evaluation
@@ VI Acknowledgements
S0.1 Molecular Geometries
Table S1: Coordinates of the methanol and 16 hydrogen dimer chain system. See Figures 5 and 6 in the paper. For shorter and longer H2 chains, the same placement pattern below is applied. The “stretched” geometry is obtained by stretching the O-H bond by 0.2 Bohr along the original bond axis. Units are in Bohr.
Z y z
C 0.088 441 23-0.088\,441\,23- 0.088 441 23 1.309 295 041.309\,295\,041.309 295 04 0.000 000 $000 0$
O 0.087 760 49-0.087\,760\,49- 0.087 760 49 1.470 368 55-1.470\,368\,55- 1.470 368 55 0.000 000 $000 0$
H 2.078 824 13-2.078\,824\,13- 2.078 824 13 1.850 932 391.850\,932\,391.850 932 39 0.000 000 $000 0$
H 0.831 955 560.831\,955\,560.831 955 56 2.081 773 802.081\,773\,802.081 773 80 1.694 536 621.694\,536\,621.694 536 62
H 0.831 955 560.831\,955\,560.831 955 56 2.081 773 802.081\,773\,802.081 773 80 1.694 536 62-1.694\,536\,62- 1.694 536 62
H 1.648 924 761.648\,924\,761.648 924 76 2.105 958 89-2.105\,958\,89- 2.105 958 89 0.000 000 $000 0$
H 0.088 441 23-0.088\,441\,23- 0.088 441 23 5.000 000 005.000\,$0 00$ 0.000 000 $000 0$
H 0.088 441 23-0.088\,441\,23- 0.088 441 23 6.393 713 296.393\,713\,296.393 713 29 0.000 000 $000 0$
H 0.588 441 23-0.588\,441\,23- 0.588 441 23 9.000 000 009.000\,$0 00$ 0.000 000 $000 0$
H 0.588 441 23-0.588\,441\,23- 0.588 441 23 10.393 713 2910.393\,713\,2910.393 713 29 0.000 000 $000 0$
H 1.088 441 23-1.088\,441\,23- 1.088 441 23 13.000 000 0013.000\,$0 00$ 0.000 000 $000 0$
H 1.088 441 23-1.088\,441\,23- 1.088 441 23 14.393 713 2914.393\,713\,2914.393 713 29 0.000 000 $000 0$
H 1.588 441 23-1.588\,441\,23- 1.588 441 23 17.000 000 0017.000\,$0 00$ 0.000 000 $000 0$
H 1.588 441 23-1.588\,441\,23- 1.588 441 23 18.393 713 2918.393\,713\,2918.393 713 29 0.000 000 $000 0$
H 2.088 441 23-2.088\,441\,23- 2.088 441 23 21.000 000 0021.000\,$0 00$ 0.000 000 $000 0$
H 2.088 441 23-2.088\,441\,23- 2.088 441 23 22.393 713 2922.393\,713\,2922.393 713 29 0.000 000 $000 0$
H 2.588 441 23-2.588\,441\,23- 2.588 441 23 25.000 000 0025.000\,$0 00$ 0.000 000 $000 0$
H 2.588 441 23-2.588\,441\,23- 2.588 441 23 26.393 713 2926.393\,713\,2926.393 713 29 0.000 000 $000 0$
H 3.088 441 23-3.088\,441\,23- 3.088 441 23 29.000 000 0029.000\,$0 00$ 0.000 000 $000 0$
H 3.088 441 23-3.088\,441\,23- 3.088 441 23 30.393 713 2930.393\,713\,2930.393 713 29 0.000 000 $000 0$
H 3.588 441 23-3.588\,441\,23- 3.588 441 23 33.000 000 0033.000\,$0 00$ 0.000 000 $000 0$
H 3.588 441 23-3.588\,441\,23- 3.588 441 23 34.393 713 2934.393\,713\,2934.393 713 29 0.000 000 $000 0$
H 4.088 441 23-4.088\,441\,23- 4.088 441 23 37.000 000 0037.000\,$0 00$ 0.000 000 $000 0$
H 4.088 441 23-4.088\,441\,23- 4.088 441 23 38.393 713 2938.393\,713\,2938.393 713 29 0.000 000 $000 0$
H 4.588 441 23-4.588\,441\,23- 4.588 441 23 41.000 000 0041.000\,$0 00$ 0.000 000 $000 0$
H 4.588 441 23-4.588\,441\,23- 4.588 441 23 42.393 713 2942.393\,713\,2942.393 713 29 0.000 000 $000 0$
H 5.088 441 23-5.088\,441\,23- 5.088 441 23 45.000 000 0045.000\,$0 00$ 0.000 000 $000 0$
H 5.088 441 23-5.088\,441\,23- 5.088 441 23 46.393 713 2946.393\,713\,2946.393 713 29 0.000 000 $000 0$
H 5.588 441 23-5.588\,441\,23- 5.588 441 23 49.000 000 0049.000\,$0 00$ 0.000 000 $000 0$
H 5.588 441 23-5.588\,441\,23- 5.588 441 23 50.393 713 2950.393\,713\,2950.393 713 29 0.000 000 $000 0$
H 6.088 441 23-6.088\,441\,23- 6.088 441 23 53.000 000 0053.000\,$0 00$ 0.000 000 $000 0$
H 6.088 441 23-6.088\,441\,23- 6.088 441 23 54.393 713 2954.393\,713\,2954.393 713 29 0.000 000 $000 0$
H 6.588 441 23-6.588\,441\,23- 6.588 441 23 57.000 000 0057.000\,$0 00$ 0.000 000 $000 0$
H 6.588 441 23-6.588\,441\,23- 6.588 441 23 58.393 713 2958.393\,713\,2958.393 713 29 0.000 000 $000 0$
H 7.088 441 23-7.088\,441\,23- 7.088 441 23 61.000 000 0061.000\,$0 00$ 0.000 000 $000 0$
H 7.088 441 23-7.088\,441\,23- 7.088 441 23 62.393 713 2962.393\,713\,2962.393 713 29 0.000 000 $000 0$
H 7.588 441 23-7.588\,441\,23- 7.588 441 23 65.000 000 0065.000\,$0 00$ 0.000 000 $000 0$
H 7.588 441 23-7.588\,441\,23- 7.588 441 23 66.393 713 2966.393\,713\,2966.393 713 29 0.000 000 $000 0$

### v2

@@ 2.6  Initialization
We begin by placing electrons according to a Lewis structure so that the molecule is roughly spatially and spin neutral. The position of each electron (except for those in the carbon and oxygen cores) is then randomized within a 1 Bohr box centered on its Lewis placement. After separating the nuclei into fragments — in our tests, each carbon and any hydrogens, oxygens, or O-H groups bonded to it form a fragment, as do individual H2 molecules — we then group the initial electron positions by fragment by assigning each electron to the closest fragment as determined by the distance to the fragment’s nuclear centroid. We burn in the electron positions using a “round-robin” approach in which we loop over each fragment and, for each one, unfreeze and sample its electrons and those of its two nearest neighbors in each direction. At the beginning of each fragment’s burn in, and once after they have all been burned in, we re-invert $M$, for a total of $n_{\mathrm{fragments}}+1$ large matrix inversions. In the systems we test, these inversions together cost much less than the initial Hartree-Fock calculation. After the round-robin burn in, we reassign electrons to fragments, again based on their proximities to the fragments’ nuclear centroids. At this point, we have our initial positions from which we separately launch the Markov chains that evaluate each fragment’s multipole and then continue on to evaluate each fragment’s energy using the approximate Hamiltonian in Eq. (8).
3  Results
3.1  Computational Details
@@
We performed all quantum chemistry calculations for this project using the PYSCF package.49, 46, 48 We optimized all geometries at the MP2/6-31G level of theory unless otherwise specified. For our Hartree-Fock references, we used restricted Slater determinants with localized Foster-Boys orbitals17 using the 6-31G(d) basis set. To ensure electron-nuclear cusps23 were satisfied for the VMC sampling, we employed a recently developed cusping algorithm40 in which all AOs are explicitly cusped at each nucleus. We used a simple two-body Jastrow factor18, 9, 8 to satisfy the electron-electron cusps (A parameter = 0.03). Regular and spotlight sampling VMC calculations were performed using our own pilot code.
@@
3.2  Impact of the Buffer Region
@@
Figure 4: Each fragment’s average uncertainty contribution to the O-H bond stretching energy in 1-octanol. Each fragment contains one C atom and any H or O-H bonded to it.
3.3  Cost Scaling
In investigating whether observed computational cost scaling matches the expectations from our analysis above, let us begin with a simple test in which we take a fixed number of samples for each fragment. As our implementation does not exploit orbital locality, we expect each sample to carry an $O(N)$ cost, which, when combined with the fact that there are $O(N)$ fragments, leads us to expect $O(N^{2})$ cost growth overall. We test this theory in a chain of hydrogen dimers (bond length = 1.393 Bohr) arranged in a line, in which each H2 molecule’s center is placed 4.03 Bohr away from the previous center, and all bond axes are arranged at an angle of 7.12 degrees relative to the line passing through the centers. Using each H2 as its own fragment and taking 110 samples per fragment, we see in Figure 3 that the spotlight sampling cost does indeed grow as $O(N^{2})$. While our current code cannot test the reduction to $O(N)$ that would result from local orbitals, that reduction is well established, 6, 15 and so we instead turn to the question of how effectively costs can be reduced by taking fewer samples in fragments farther away from the local chemical change in question.
Figure 5: Each fragment’s average uncertainty contribution to the O-H bond stretching energy in methanol-$\mathrm{(H_{2})_{16}}$. Each molecule is one fragment.
To address this question, we first perform tests of how much uncertainty each fragment contributes to a bond stretching energy when we take a fixed number of samples per fragment. We perform our first test in 1-octanol, where we perform 150 runs on each fragment, in which each run uses 720,000 samples. For each run in each fragment, we use a blocking analysis to evaluate the uncertainty of that fragment’s contribution to the bond stretching energy in that run. The averages of these uncertainties over all 150 runs are plotted in Figure 4, where we see that the uncertainties decrease as one moves away from the CH2OH fragment (fragment 1) that contains the bond stretch. In fact, these uncertainties decay fast enough that they are bounded by an $a/l^{2}$ function, with $a=0.01$ Hartree and $l$ the fragment number. As discussed in Section 2.1, this is a highly desirable bound, as it suggests that decreasing the sample size for farther fragments should be highly effective.
@@ 3.3  Cost Scaling
Seeing that uncertainties decay rapidly as one gets farther from the local chemical change, we now exploit this behavior to demonstrate that, even without making use of local orbital accelerations, spotlight sampling can achieve lower costs than traditional VMC. To do so, we again calculate the O-H bond stretching energy for methanol-$\mathrm{(H_{2})_{n}}$, but this time we employ blocks of 960,000 samples for the first three fragments and much smaller blocks of 4,000 samples for the remaining fragments, versus whole-system blocks of 960,000 samples for standard VMC. In each system, we compute enough sampling blocks so that the overall uncertainty in the stretching energy is below 0.0005 Hartree, always using at least 15 blocks to ensure a reliable uncertainty estimate.
@@
In Figure 6, we see that this test of spotlight sampling achieves near-linear scaling with system size, with spotlight’s cost lower than standard correlated sampling once there are about 100 electrons. The linear scaling is expected, as the vast majority of the samples are in the first three fragments, making the total number of samples nearly independent of system size. Indeed, the linear cost growth comes instead from our code’s $O(N)$ cost per sample. The small quadratic term in the cost growth that begins to be visible in the spotlight timing data in the largest system is due to the fact that we used fixed-size 4,000 sample blocks even for the farthest fragments, and could be eliminated by tapering this sample size further. It is important to re-emphasize that this crossover with standard sampling has been achieved without exploiting local orbitals. Were one to do so, the crossover would occur even earlier, and the cost growth of spotlight sampling could drop to sub-linear, assuming local Jastrows and fast multipole methods were also employed. Although each spotlight prediction agreed with standard sampling within error bars, this system is somewhat artificial, and so we turn to a series of alcohols and systems with $\pi$ conjugation as more realistic tests of the method’s accuracy in Section 3.5.
3.4  Statistical Efficiency Comparison
We have also tested whether the width of the Gaussian from which one-electron moves are drawn (which is the analogue of the diffusion Monte Carlo timestep) has an effect on statistical efficiency, and whether the effect is different in standard versus spotlight sampling. Specifically, we have performed six evaluations of the 1-octanol stretching energy, three each for standard and spotlight sampling. Across the three tests, we varied the standard deviation of the Gaussian from which one-electron moves were drawn, testing values of 0.3, 0.6 and 1.0 Bohr. As each test used the same total number of samples, we take the uncertainty in the stretching energy as our measure of statistical efficiency. Figure 7 shows that neither standard nor spotlight sampling had its statistical efficiency much affected by the one-electron draw width and that the spotlight approach is slightly more efficient by this measure. As seen in Table 1, the acceptance ratios for standard and spotlight sampling are almost identical and display the expected trend of decreasing as the draw width increases. The modest difference in statistical efficiency for the two methods is instead more likely caused by the fact that long-range electrostatics are handled by multipole expansions in the spotlight approach, which introduce different amounts of noise into the long-range electron-electron repulsion energy than standard VMC’s approach of averaging repulsions between explicit electron positions.
Figure 7: The uncertainty of the O-H bond stretch energy in 1-octanol as a function of the standard deviation of the Gaussian distribution for the 1-electron move proposals.
Table 1: Comparison of the 1-electron move acceptance ratio for different standard deviations of the move proposal distribution. For the spotlight sampling scheme, the acceptance ratio is taken as the average ratio over all fragments.
Standard deviation [a.u.]: 0.3 0.6 1.0
Accept ratio (standard) 0.61 0.42 0.27
Accept ratio (spotlight) 0.61 0.42 0.28
3.5  Accuracy
3.5.1 Alcohol Chains
@@
To test whether our approximations to the Hamiltonian and probability distribution produce meaningful errors in predicted energy differences, we apply the spotlight sampling approach to predict the energetic cost of a 0.2 Bohr O-H bond stretch in a series of alcohols: 1-butanol through 1-octanol. As seen in Figure 8, spotlight sampling reproduces the predictions of standard correlated sampling in all cases. We note that this accuracy comes despite the fact that, in octanol, the Markov chain that estimates the contribution from the fragment containing the stretched bond froze more than half of the electrons.
@@
Figure 8: Predictions for the energy needed to stretch an O-H bond in the alcohols 1-butanol through 1-octanol.
3.5.2 Conjugated Molecules
Figure 9: Predictions for the energy needed for a 0.2 Bohr stretch of the central C=C bond in CCO(CH2CHCHCH3)2 as well as a depiction of the molecule and the zoning when the spotlight is placed on the central CCO group.
To test spotlight sampling’s accuracy in settings where the chemical bonding is less local in character, we have also explored double-bond stretching in a pair of molecules with differing degrees of $\pi$-system delocalization. In the CCO(CH2CHCHCH3)2 molecule shown in Figure 9, the four different double bonds are separated by CH2 groups into three distinct $\pi$ systems, and the simple zoning scheme tested in this study leads spotlight to nicely reproduce the results of standard VMC. This success comes despite the fact that we are now stretching a double bond and despite the fact that, when that bond is in the spotlight A region, the other carbon-carbon double bonds are each split between the C and D regions, meaning that some of their electrons are frozen while others are moving.
In contrast, the CCO(CHCHCH3)2 molecule shown in Figure 10 has all four of its double bonds linked together in one delocalized $\pi$ system, and in this case the spotlight approach using our simple zoning scheme is less accurate. As the $\pi$ molecular orbitals now extend across most of the molecule, the sampling of electrons in the stretched CCO group when that group is in the spotlight (the zoning for which is shown in Figure 10) is now significantly affected by the fact that the electrons in the methyl caps at either end are frozen. Although they are far away from the CCO group in terms of distance through space, the delocalized $\pi$ system extends to the edge of those methyl groups and so experiences spurious Pauli exclusion effects from the frozen electrons. As it is one of the bonds in this $\pi$ system that is being stretched, it is perhaps not surprising that accuracy suffers.
Figure 10: Predictions for the energy needed for a 0.2 Bohr stretch of the central C=C bond in CCO(CHCHCH3)2 as well as a depiction of the molecule and the zoning when the spotlight is placed on the central CCO group.
Seeing these two related double-bond stretches lead the spotlight approach to one accurate and one inaccurate outcome raises questions about how best to generalize and automate the zoning setup in the future. In particular, these results suggest that it may be important to iterate the spotlight scheme by asking the following question: does the simple zoning approach used here give the same prediction as one in which the key zones nearest the local chemical change are fused together into a larger zone? If not, further expansion of this central zone can be tested until the prediction stops changing. While this iteration may be unnecessary when delocalized $\pi$ systems are not present, it would provide an automated way to discover the extent of the nonlocal behavior, at least insofar as it affects the spotlight sampling prediction. In addition to iteration, it will likely also be important to have two criteria for choosing which zones should be B zones for a given A zone: the current one based on molecular connectivity but also a second one based on distance. In larger molecules that can fold over on themselves, bonding connectivity may be a poor guide to which other zones will have the strongest effects on the currently-spotlighted A zone.
4  Conclusion
@@
This study has explored the efficacy of a spotlight sampling approach to predicting the energy differences associated with local chemical changes in VMC. In this approach, standard correlated sampling is approximated by using a series of separate Markov chains, one for each fragment in the molecule, in which only the electrons near that fragment move. Our theoretical analysis suggests that the cost of the approach can be made linear with system size when employing nonlocal orbitals, and that this cost should become sub-linear when local orbitals, local Jastrows, and fast multipole methods are used. Timing calculations show that the overall cost grows quadratically when we take a fixed number of samples per spotlighted fragment, and that this cost growth becomes essentially linear, as expected, once we exploit the fact that per-fragment uncertainty decays rapidly for fragments farther from the local change. To maintain accuracy despite the biases inherent to freezing most electrons’ positions, we have developed an approximate fragment Hamiltonian that employs multipoles for long range interactions. Tests on a series of alcohols and molecules with $\pi$-conjugation show that this approach can reproduce the predictions of standard sampling within statistical uncertainty.
In future work, there are important questions to answer about numerical stability and the applicability of the approach to more sophisticated wave functions. It appears highly likely that simply employing pseudopotentials will go a long way towards mitigating numeric challenges, but this expectation needs to be established firmly with data. Regarding other wave functions, a generalization of the current scheme should be applicable to multi-determinant wave functions thanks to the fact that, in practice, such wave functions are evaluated using an extended Slater matrix that will have the same pattern of rows that do and do not change when electrons in only one region are moving. In contrast, for neural network and backflow wave functions in which moving one electron can change every element of the Slater matrix, it may be necessary to impose some “row locality” on the network setup in order for spotlight sampling to be helpful.
5  Appendix
5.1  Correlated Sampling with the Space-Warp Coordinate Transformation
@@
We employ the correlated sampling 29 technique known as the space-warp coordinate transformation 51 to reduce the statistical uncertainty of simple bond stretches (e.g., O-H stretch in 1-n-ol, n = butan-, pentan-, hexan-,…). The relative energy difference for such a stretch, $E_{s}-E$, is rewritten to be sampled from the unstretched $\Psi$ reference:
@@
$J(R_{i})$ is the Jacobian of the space-warp coordinate transformation, 51, 16 and
@@
5.2  Kinetic Energy Evaluation